A rectangles length is 3 times its width. The area is 27 square inches. Find the perimeter.
Answer by JulietG(1812)

L = 3W
A = L * W
27 = 3W * w
27 = 3W^2
Divide each side by 3
9 = w^2
3 = W
If the width is 3, then the length is 9
Perimeter = 2L + 2W
P = (2*9) + (2*3)
P = 18 + 6
P = 24